OpenStudy (anonymous):

Find dy/dx. y=(3x^4+4x^3-1)/x^3

OpenStudy (anonymous):

\[y = \frac{ 3x^4+4x^3-1 }{ x^3 }\]

OpenStudy (anonymous):

So what rule would you use? :)

OpenStudy (anonymous):

quotient rule, right?

OpenStudy (psymon):

Break it into 3 fractions o.o

OpenStudy (psymon):

\[\frac{ 3x^{4}+4x^{3}-1 }{ x^{3} }= \frac{ 3x^{4} }{ x^{3} }+\frac{ 4x^{3} }{ x^{3} }-\frac{ 1 }{ x^{3} }\]

OpenStudy (anonymous):

originally, I got 3x^6+3x^2/x^5 but I don't know how to fix it to make it correct

OpenStudy (psymon):

Okay, well you can do it with the 3 fractions, because that ends up making it much easier. But is that whatyou got trying to do quotient rule?

OpenStudy (anonymous):

yeah, that's what I was taught but i did something wrong

OpenStudy (psymon):

Well, I mean, of course you cna quotient rule, I just mean to say that there's an easier way with this problem that would be nice to try to recognize. But okay, lets say we HAD to do it with quotient rule or get an F or something: \[\frac{ f'(x)g(x) - f(x)g'(x) }{ [g(x)]^{2} }\] SO f is the top function and g is the bottom function. Ill just stack the information and then combine it in later: f(x) = 3x^4 +4x^3 - 1 g(x) = x^3 f'(x) = 12x^3 + 12x^2 g'(x) = 3x^2 \[\frac{ (12x^{3}+12x^{2})(x^{3}) - (3x^{4}+4x^{3}-1)(3x^{2}) }{ x^{6} }\] \[\frac{ 12x^{6}+12x^{5}-(9x^{6}+12x^{5}-3x^{2}) }{ x^{6} }\] \[\frac{ 12x^{6} + 12x^{5} - 9x^{6} - 12x^{5} + 3x^{2} }{ x^{6} }\] \[\frac{ 3x^{6}+3x^{2} }{ x^{6} } = \frac{ 3x^{2}(x^{4}+1) }{ x^{6} }\implies \frac{ 3(x^{4}+1) }{ x^{4} }\] And you can separate that into two fractions if you wanted,but yeah.

OpenStudy (anonymous):

okay, i got as far as the last line there until it simplified into the final answer. how did you get x^4 in the denominator?

OpenStudy (psymon):

Can you see how I factored it into 3x^2(x^4+1) on top? do you understand up to that point or is it further back?

OpenStudy (anonymous):

I understand right up to that point

OpenStudy (anonymous):

ohhh do i subract the exponents? is that how the x^4 got in the denominator?

OpenStudy (psymon):

Well, I just factor the 3x^6 + 3x^2 portion. In the same way I could factor something like 3x+9 into 3(x+3), I just did that with the top. That allows me to pull out a greatest factor of 3x^2 to give me 3x^2(x^4+1). And yes, exactly. Or you could even think of it like this: